A car’s tachometer indicates the angular velocity of the engine’s crankshaft in rpm. A car stopped


at a traffic light has its engine idling at 500 rpm. When the light turns green, the crankshaft


rotation speeds up at a constant rate to 2500 rpm in a time interval of 3.0 s. How many


revolutions does the crankshaft make in these time interval?

1
Expert's answer
2022-02-11T08:46:05-0500

Given:

"n_1=500\:\rm rpm"

"n_2=2500\:\rm rpm"

"t=3.0\:\rm s"


The number of revolution

"N=\frac{n_1+n_2}{2}t=\frac{500+2500}{2}*3.0=4500"